OpenStudy (anonymous):

Complete the inverse of F(x) below. Do not include decimals or "F-1(y) =" F(x) = 2x - 2 F-1(y) = _??

OpenStudy (lukecrayonz):

To find the inverse, you replace y with x in the equation. I'm going to assume you know F(x) is other words for y.

OpenStudy (lukecrayonz):

So y=2x-2, replace y for x, and x for y.

OpenStudy (lukecrayonz):

x=2y-2

OpenStudy (lukecrayonz):

Solve:)

OpenStudy (lukecrayonz):

x+2=2y y=(x+2)/2

OpenStudy (anonymous):

how would you solve this? This just confuses me so much

OpenStudy (lukecrayonz):

Let's say the equation is F(x)=5x+10 Replace x with y, and y with x. F(x) is just other words for y, so, x=5y+10 x=5y+10 -10 x-10=5y (x-10)/5=y F-^1(y)=(x-10)/5

OpenStudy (anonymous):

2y+2/2 ?